# Global Bayesian Analysis of $\mathrm{J}/\psi$ Photoproduction on Proton and Lead Targets

### v1

@@ I Introduction
This tension has not been resolved to date, motivating the current work. [^1] Recently, it has been argued in Ref. [31] that including a subset of NLO corrections may resolve this issue, but this conclusion is not confirmed by a full NLO calculation [32]. Here, we set out to perform a combined Bayesian analysis of both $\gamma+p$ and $\gamma+\mathrm{Pb}$ diffractive $\mathrm{J}/\psi$ production data, both coherent and incoherent, to determine whether one can find any set of parameter values that allows for the simultaneous description of both systems.
@@ II Exclusive vector meson production at high energy
×$\times\left\{m_{c}^{2}K_{0}(\varepsilon r)\phi_{\rm T}(r,z)\right.$
@@
We also consider a modification to the MV model that suppresses the high-frequency modes (scattering amplitude for small dipoles) following Ref. [24]. This is motivated by the fact that leading-order dipole model fits to structure function data prefer dipole-proton scattering amplitudes that decrease faster than the MV model prediction $\sim r^{2}$ as $r$ goes to 00 [45, 46]. In practice, this is achieved by introducing a ultraviolet (UV) damping factor $v_{\mathrm{UV}}$ as
@@ III Bayesian inference setup
×$~{}~{}~{}~{}\times\exp\left[-\frac{1}{2}(y({\boldsymbol{\theta}})-y_{\mathrm{exp}})^{\mathsf{T}}\Sigma^{-1}(y({\boldsymbol{\theta}})-y_{\mathrm{exp}})\right].$ (18)
@@ IV.1 Standard setup
(a) $\gamma+p\to\mathrm{J}/\psi+p$ cross section as a function of center-of-mass energy compared to ALICE [78, 79], H1 [75, 76], ZEUS [77], and LHCb [80] data.
@@ IV.2 Model extensions
Model $A$ Model $B$ $\ln(\mathcal{B}_{A/B})$
@@
standard $\omega,N_{q}$ $-2.18\pm 0.02$
standard $N_{q}$ $-3.12\pm 0.02$
standard $\omega,v_{\rm UV},N_{q}$ $-1.03\pm 0.02$
standard $K$ $-18.56\pm 0.02$
@@ IV.3 Global analysis with and without a $K$ factor
(a) $\gamma+p\to\mathrm{J}/\psi+p$ cross section as a function of center-of-mass energy compared to ALICE [78, 79], H1 [75, 76], ZEUS [77], and LHCb [80] data.

### v2

@@ I Introduction
This tension has not been resolved to date, motivating the current work. [^1] Recently, it has been argued in Ref. [31] that including a subset of next-to-leading order (NLO) corrections may resolve this issue, but this conclusion is not confirmed by a full NLO calculation [32]. Here, we set out to perform a combined Bayesian analysis of both $\gamma+p$ and $\gamma+\mathrm{Pb}$ diffractive $\mathrm{J}/\psi$ production data, both coherent and incoherent, to determine whether one can find any set of parameter values that allows for the simultaneous description of both systems.
@@ II Exclusive vector meson production at high energy
$\times\left\{m_{c}^{2}K_{0}(\varepsilon r)\phi_{\rm T}(r,z)\right$.
@@
We also consider a modification to the MV model that suppresses the high-frequency modes (scattering amplitude for small dipoles) following Ref. [24]. This is motivated by the fact that leading-order dipole model fits to structure function data prefer dipole-proton scattering amplitudes that decrease faster than the MV model prediction $\sim r^{2}$ as $r$ goes to 0 [45, 46]. In practice, this is achieved by introducing a ultraviolet (UV) damping factor $v_{\mathrm{UV}}$ as
@@ III Bayesian inference setup
$~~~~\times\exp\left[-\frac{1}{2}(y({\boldsymbol{\theta}})-y_{\mathrm{exp}})^{\mathsf{T}}\Sigma^{-1}(y({\boldsymbol{\theta}})-y_{\mathrm{exp}})\right].$ (18)
@@ IV.1 Standard setup
(a) $\gamma+p\to\mathrm{J}/\psi+p$ cross section as a function of center-of-mass energy compared to ALICE [78, 79], H1 [75, 76], ZEUS [77], and LHCb [80, 87] data. Datapoints with open markers are not included in the fit.
@@ IV.2 Model extensions
Model AA Model $B$ $\ln(\mathcal{B}_{A/B})$
@@
standard $\omega,N_{q}$ $2.18\pm 0.02$
standard $N_{q}$ $3.12\pm 0.02$
standard $\omega,v_{\rm UV},N_{q}$ $1.03\pm 0.02$
standard $K$ $18.56\pm 0.02$
@@ IV.3 Global analysis with and without a $K$ factor
(a) $\gamma+p\to\mathrm{J}/\psi+p$ cross section as a function of center-of-mass energy compared to ALICE [78, 79], H1 [75, 76], ZEUS [77], and LHCb [80, 87] data. Datapoints with open markers are not included in the fit.
@@ V Conclusions
This indicates that calculations performed entirely in the linear regime should not be able to describe the $\gamma+p$ and $\gamma+{\rm Pb}$ data simultaneously, as was also shown in a comparison between calculations using the non-linear Balitsky–Kovchegov (BK) and linear Balitsky–Fadin–Kuraev–Lipatov (BFKL) equations [90, 91, 92] in [93]. NLO BFKL calculations find an improvement, but tension with the $\gamma+{\rm Pb}$ data remains [94] (also see [95] for a related study in $\gamma+p$). A future Bayesian analysis utilizing BFKL instead of JIMWLK evolution will help quantify the need for non-linear effects. Calculations of exclusive vector meson production in $\gamma+{\rm Pb}$ collisions using nuclear parton distribution functions (nPDFs) [96] within the collinear factorization framework [97] show good agreement with experimental data, yet scale and nPDF uncertainties are large. Resumming higher-order corrections can significantly suppress the large scale uncertainty, which has been done for $\gamma+p$ collisions in [98] using high-energy factorization [99, 100, 101, 102], but to our knowledge is still missing for the $\gamma+$Pb case.